A chemist has 200 milliliters of a solution that is 20% acid. She needs a solution that is 5% acid. How many milliliters of water should she add?
Answer by robertb(5830)

200*0.20 + x*0.00 = (200 + x)*0.05
==> 40 = 10 + 0.05x
==> 30 = 0.05x
==> x = 30/0.05 = 600 milliliters of water should be added.